# Position Paper: Challenges and Opportunities in Topological Deep Learning

### v1

@@ 6.2 Topological Representation Learning
There are several possible interpretations of the umbrella term ‘topological representation learning’. A first interpretation would be to develop training strategies that enable a neural network to elicit higher-order information from the data for the task at hand. An immediate follow-up question in this context is whether this uncovered higher-order structure is meaningful in any semantically valid sense. In the context of GNNs, it is, for example, unclear whether such an uncovered higher-order structure might not be equivalent to the message-propagation phase in the spirit of Veličković (2022). However, even if this is true, it might still be possible to analyze these structures to improve the explainability and transparency of a neural network.
A second interpretation of topological representation learning would be to develop methods that promote or penalize certain topological properties of a neural network’s internal representation of the data or its output. On the one hand, such properties could be informed by a priori knowledge or corresponding properties in the input domain, as successfully demonstrated in image segmentation problems (Hu et al., 2021b; Gupta et al., 2023), learning representations to synthesize new shapes (Waibel et al., 2022), using random walks on complexes to obtain cell embeddings (Billings et al., 2019; Schaub et al., 2020), or preserving topological properties while training autoencoders (Moor et al., 2020; Hajij et al., 2020). On the other hand, topological characteristics of representations could also be guided by establishing provably beneficial properties for generalization, as proposed in Hofer et al. (2019, 2020a). Another facet of this interpretation of representation learning involves the analysis and modification of the decision boundaries of classifiers. Here, Chen et al. (2019) have shown that regularizers based on persistent homology lead to better generalization and higher predictive performance.
@@
The explainability of predictions made by GNNs (Luo et al., 2020; Lucic et al., 2022) is complicated by the fact that even a small subgraph is difficult for a human to visualize and grasp quickly. Such complications are amplified in TDL, since visualization and human apprehension of higher-order domains is virtually nonexistent to date. Nevertheless, information in a higher-order domain can be summarized more flexibly by a smaller set of cells or hyperedges, leading to potentially interpretable structure. In other words, the richer structure attained by higher-order relations may enable more effective representation learning. For example, the work of Battiloro et al. (2023b) on latent topological inference (LTI) has introduced a trainable method to identify higher-order interactions between entities, combining advances in TDL and differentiable sampling. Building upon LTI might lead to a general framework for topological representation learning on arbitrary higher-order domains. Building on the graph-based representation learning approach of Hamilton et al. (2017), the work of Hajij et al. (2022b, 2023b) has introduced a topological representation learning framework for neighbourhood-based methods, which provides scope for topological neural network architecture search.
@@
Higher-order relations, encoded by cells or hyperdeges, offer candidate components for effective topological representation learning. Latent topological inference outlines a tentative methodological path toward generalized representation learning for TDL. Existing work on topological representation learning provides a means to develop topological neural network architecture search algorithms.
@@ 7 Final Remarks
TDL offers immense potential to transform various domains. However, to truly harness this potential and accelerate the transition of TDL from theoretical constructs to real-world applications, the research community is invited to adopt a multidisciplinary approach rooted in innovation, collaboration, and open science. Such an approach necessitates the convergence of diverse fields, bringing together experts in mathematics, computer science, and machine learning, to foster a rich cross-pollinating environment. By encouraging open dialogue and the exchange of ideas, TDL can evolve rapidly, breaking through current barriers, and sparking groundbreaking advancements.
@@ Acknowledgements
This work is partially supported by DOE grant DE-SC0023157 (BW), NIH grants R01AI164266 (GWW) and R35GM148196 (GWW), and NSF grants 2134241 (NM, MH), DMS-2134223 (BW), IIS-2205418 (BW), DMS2052983 (GWW) and IIS-1900473 (GWW). This work was partially funded by the Italian NRRP (PE00000001 - program ‘RESTART’), and by the 6G-GOALS project under the 6G SNS-JU Horizon program, n.101139232 (PDL). VM is partially funded by NSF DMS No 2012609, US Army Research Lab Contract No W911NF-21-2-0186, US Army Research Lab Contract No W911NF-22-2-0143, US Army Research Office No W911NF-21-1-0094, and The University of Tennessee Materials Research Science & Engineering Center – The Center for Advanced Materials and Manufacturing – NSF DMR No. 2309083. BR is supported by the Bavarian state government with funds from the Hightech Agenda Bavaria. SS is partly funded by the Sapienza grant RM1221816BD028D6 (DESMOS). MTS acknowledges funding by the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
The authors thank Cristian Bodnar for his input on the theoretical foundations of TDL, and Thomas Kipf and Charles Blundell for their reviews and feedback.
@@ A.1 Higher-Order Deep Learning Models
Recent years have witnessed a growing interest in higher-order networks (Mendel, 1991; Battiston et al., 2020; Bick et al., 2021) due to their effective ability to capture higher-order interactions. In signal processing and deep learning, various approaches, such as Hodge-theoretic methods, message-passing schemes, and skip connections, have been developed for higher-order deep learning models.
@@
In the context of message passing on complexes, a general message-passing scheme on topological spaces has been developed by Hajij et al. (2020), encompassing previous proposals (Gilmer et al., 2017; Ebli et al., 2020; Bunch et al., 2020; Hayhoe et al., 2022) and employing various topological neighborhood aggregation schemes. The expressive power of simplicial and cellular message passing neural networks has been studied by Bodnar (2022) and Bodnar et al. (2022), respectively. The work of Hajij et al. (2023b) provides the mathematical blueprint of topological deep learning, unifying existing deep learning architectures with a common mathematical language. Papillon et al. (2023) has conducted a review of message-passing topological neural network architectures, employing foundational concepts from Hajij et al. (2023b). This survey highlights key challenges and promising directions for future research in the field. Recurrent simplicial neural networks for trajectory prediction have been introduced (Mitchell et al., 2023), and a multi-signal approach on higher-order networks utilizing the Dirac operator has been developed (Calmon et al., 2022; Hajij et al., 2023a).

### v3

@@ 6.2 Topological Representation Learning
There are several possible interpretations of the umbrella term TRL. As a high-level definition, TRL is concerned with automatically learning and exploiting multiscale topological descriptions of the data and of the neural network’s internal embeddings, which scale efficiently to larger dimensions and which provide benefits in efficiency and transparency compared to the equivalent graph-based representations.
One question in TRL is how to develop training strategies that enable a neural network to elicit higher-order information from the data for the task at hand. An immediate follow-up question in this context is whether this uncovered higher-order structure is meaningful in any semantically valid sense. In the context of GNNs, it is, for example, unclear whether such an uncovered higher-order structure might not be equivalent to the message-propagation phase in the spirit of Veličković (2022). However, even if this is true, it might still be possible to analyze these structures to improve the explainability and transparency of a neural network.
Another question in TRL is how to develop methods that promote or penalize certain topological properties of a neural network’s internal representation of the data or its output. On the one hand, such properties could be informed by a priori knowledge or corresponding properties in the input domain, as successfully demonstrated in image segmentation problems (Hu et al., 2021b; Gupta et al., 2023), learning representations to synthesize new shapes (Waibel et al., 2022), using random walks on complexes to obtain cell embeddings (Billings et al., 2019; Schaub et al., 2020), or preserving topological properties while training autoencoders (Moor et al., 2020; Hajij et al., 2020; Trofimov et al., 2023). On the other hand, topological characteristics of representations can be guided by establishing provably beneficial properties for generalization (Hofer et al., 2019, 2020a) or can be used for comparing neural network representations (Barannikov et al., 2022). Another facet of this interpretation of representation learning involves the analysis and modification of the decision boundaries of classifiers. Here, Chen et al. (2019) have shown that regularizers based on persistent homology lead to better generalization and higher predictive performance.
@@
The explainability of predictions made by GNNs (Luo et al., 2020; Lucic et al., 2022) is complicated by the fact that even a small subgraph is difficult for a human to visualize and grasp quickly. Such complications are amplified in TDL, since visualization and human apprehension of higher-order domains is virtually nonexistent to date. Nevertheless, information in a higher-order domain can be summarized more flexibly by a smaller set of cells or hyperedges, leading to potentially interpretable structure. In other words, the richer structure attained by higher-order relations may enable more effective representation learning. For example, the work of Battiloro et al. (2023b) on latent topological inference (LTI) has introduced a trainable method to identify higher-order interactions between entities, combining advances in TDL and differentiable sampling. Building upon LTI might lead to a general framework for TRL on arbitrary higher-order domains. Building on the graph-based representation learning approach of Hamilton et al. (2017), the work of Hajij et al. (2022b, 2023b) has introduced a TRL framework for neighbourhood-based methods, which provides scope for topological neural network architecture search.
@@
Higher-order relations, encoded by cells or hyperdeges, offer candidate components for effective TRL. Latent topological inference outlines a tentative methodological path toward generalized representation learning for TDL. Existing work on TRL provides a means to develop topological neural network architecture search algorithms.
@@ 7 Final Remarks
To accelerate the transition of TDL from theoretical constructs to real-world applications, the research community is invited to adopt a multidisciplinary approach rooted in innovation, collaboration, and open science. Such an approach necessitates the convergence of diverse fields, bringing together experts in mathematics, computer science, and machine learning, to foster a rich cross-pollinating environment. By encouraging open dialogue and the exchange of ideas, TDL can evolve rapidly, breaking through current barriers, and sparking groundbreaking advancements.
@@ Acknowledgements
This work is partially supported by DOE grant DE-SC0023157 (BW), NIH grants R01AI164266 (GWW) and R35GM148196 (GWW), and NSF grants 2134241 (NM, MH), DMS-2134223 (BW), IIS-2205418 (BW), CCF-2112665 (YW), CCF-2310411 (YW), DMS2052983 (GWW) and IIS-1900473 (GWW). This work was partially funded by the Italian NRRP (PE00000001 - program ‘RESTART’), and by the 6G-GOALS project under the 6G SNS-JU Horizon program, n.101139232 (PDL). VM is partially funded by NSF DMS No 2012609, US Army Research Lab Contract No W911NF-21-2-0186, US Army Research Lab Contract No W911NF-22-2-0143, US Army Research Office No W911NF-21-1-0094, and The University of Tennessee Materials Research Science & Engineering Center – The Center for Advanced Materials and Manufacturing – NSF DMR No. 2309083. BR is supported by the Bavarian state government with funds from the Hightech Agenda Bavaria. SS is partly funded by the Sapienza grant RM1221816BD028D6 (DESMOS). MTS acknowledges funding by the European Union (ERC, HIGH-HOPeS, 101039827). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
The authors thank Cristian Bodnar for his input on the theoretical foundations of TDL, and the ICML reviewers, Thomas Kipf and Charles Blundell for their reviews and feedback.
@@ A.1 Higher-Order Deep Learning Models
Recent years have witnessed a growing interest in higher-order networks (Mendel, 1991; Battiston et al., 2020; Bick et al., 2023) due to their effective ability to capture higher-order interactions. In signal processing and deep learning, various approaches, such as Hodge-theoretic methods, message-passing schemes, and skip connections, have been developed for higher-order deep learning models.
@@
In the context of message passing on complexes, a general message-passing scheme on topological spaces has been developed by Hajij et al. (2020), encompassing previous proposals (Gilmer et al., 2017; Ebli et al., 2020; Bunch et al., 2020; Hayhoe et al., 2022) and employing various topological neighborhood aggregation schemes. The expressive power of simplicial and cellular message passing neural networks has been studied by Bodnar (2022) and Bodnar et al. (2022), respectively. The work of Hajij et al. (2023b) provides the mathematical blueprint of topological deep learning, unifying existing deep learning architectures with a common mathematical language. Papillon et al. (2023b) has conducted a review of message-passing topological neural network architectures, employing foundational concepts from Hajij et al. (2023b). This survey highlights key challenges and promising directions for future research in the field.  Zia et al. (2024) have carried out a review of TDL, focusing primarily on how the use of TDA techniques has evolved over time to support deep learning frameworks. Recurrent simplicial neural networks for trajectory prediction have been introduced (Mitchell et al., 2024), and a multi-signal approach on higher-order networks utilizing the Dirac operator has been developed (Calmon et al., 2022; Hajij et al., 2023a).